OpenStudy (anonymous):

Find the derivative and simplify dy/dx if y(x)=square root x^2 +6 Sorry, I am not sure how to make a square root symbol over the x^2+6 on here.

OpenStudy (anonymous):

Let \(f(x)=\sqrt{x^2+6}\). Then\[f'(x)=\frac{2x}{2\sqrt{x^2+6}}.\]

OpenStudy (anonymous):

Simplify it.

OpenStudy (anonymous):

Are you interested in knowing how I got that or did you just want an answer?

OpenStudy (anonymous):

x/squareroot x^2+6 this would be simplified right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

just a question - how can it be simplified with radical in the denominator?

OpenStudy (anonymous):

It's called "rationalizing" the function. Basically,\[f'(x)=\frac{x}{\sqrt{x^2+6}}=\frac{x\sqrt{x^2+6}}{x^2+6}.\]

OpenStudy (anonymous):

I thought your answer was going to leave the sq rt in the denominator. I understand rationalizing and was just ensuring that in order to be considered simplified that you still had to do that.

OpenStudy (anonymous):

Some people don't consider rationalizing the function as simplifying it.